OpenStudy (anonymous):

Need help with 2 algebra questions, will give best response if you answer correctly

OpenStudy (anonymous):

do you knovv hovv the vvhole subtract vvhen dividing thing vvorks?

OpenStudy (anonymous):

i can't do them for you. sorry.

OpenStudy (anonymous):

you don't understand how to do them?

OpenStudy (anonymous):

i do understand but its going to take a decent amount of time to explain this to you.

OpenStudy (anonymous):

can you tell me if my best guesses on them are right or not at least?

OpenStudy (anonymous):

i think its either A or B for #4 & it's either C or D for #5.

OpenStudy (anonymous):

i can tell you this. for number 4 it is not c and 5 it is not d

OpenStudy (anonymous):

hold on give me a minute that is vvrong.

OpenStudy (anonymous):

may have given you the ansvvers but dang

OpenStudy (anonymous):

im so confused i wish 2 problems didnt have to be so difficult.

jigglypuff314 (jigglypuff314):

you know how when you subtract fractions they have to have the same denominator? \[\frac{ 1 }{ 2 } + \frac{ 1 }{ 3 } \rightarrow \frac{ 1 }{ 2 }*\frac{ 3 }{ 3 } + \frac{ 1 }{ 3 }*\frac{ 2 }{ 2 } \rightarrow \frac{ 3 }{ 6 } + \frac{ 2 }{ 6 } \rightarrow \frac{ 5 }{ 6 }\] do you understand this? question 4 is the same idea...

OpenStudy (anonymous):

yes I see that

jigglypuff314 (jigglypuff314):

so \[\frac{ 3x }{ (x+1) } - \frac{ (x-5) }{ (x+6) } \rightarrow \frac{ 3x }{ (x+1) }*\frac{ (x+6) }{ (x+6) } + \frac{ (x-5) }{ (x+6) }*\frac{ (x+1) }{ (x+1) }\]

OpenStudy (anonymous):

so #4 would be C?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

not quite... you would simplify the top 3x(x+6) - (x-5)(x+1) 3x^2 + 18x - (x^2 - 4x - 5) then simplify more. what do you get?

OpenStudy (anonymous):

if I add 18x and 4x it would be B but if I'm supposed to subtract those it would be A.

OpenStudy (anonymous):

I feel like I'm supposed to add them because it's negative so it's B?

jigglypuff314 (jigglypuff314):

3x^2 + 18x - (x^2 - 4x - 5) you distribute the negative 3x^2 + 18x - x^2 + 4x + 5 so B is correct :)

OpenStudy (anonymous):

yay! now all I need is the final problem which is 5.

OpenStudy (anonymous):

with this one I add 10b and 5b right?

OpenStudy (anonymous):

because if so it must be D.

jigglypuff314 (jigglypuff314):

umm, I'm not quite sure what you mean, it was the same idea as the last one so \[\frac{ 10b }{ (b-2)(3b+1) } + \frac{ (5b+1) }{ (3b + 1) } = \frac{ 10b }{ (b-2)(3b+1) } + \frac{ (5b+1) }{ (3b + 1) }*\frac{ (b-2) }{ (b-2) }\]

jigglypuff314 (jigglypuff314):

you can't add fractions that don't have the same denominator, remember that :)

OpenStudy (anonymous):

ohh okay. so it would either be A or B then. I'm going to think .... it's A?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

yay! thank you so much for being patient with me. you've helped a lot. :) gave you best response btw.

jigglypuff314 (jigglypuff314):

glad I could help :D